# Soft Theorems and Spontaneous Symmetry Breaking

### v2

@@ Abstract
The soft photon and soft graviton theorems of Weinberg are known to derive from conservation laws associated with asymptotic symmetries. Within the corresponding classical theories, one often speaks of spontaneous symmetry breaking and vacuum degeneracy, but a genuine quantum description of this phenomenon has largely been lacking. Here we establish spontaneous symmetry breaking of asymptotic symmetries and the existence of Goldstone ‘particles’ using exclusively the language of quantum field theory. This is made possible through the reformulation of massless scattering theory in terms of carrollian conformal field theory, and the observation that soft theorems correspond to Ward identities of broken symmetries. A suitable version of Goldstone theorem shows that there must exist zero-momentum particles described by conformal fields on the celestial sphere, in agreement with the common lore. More specifically, these belong to unitary representations in the discrete series of the Lorentz group, and are therefore naturally equipped with logarithmic two-point functions. We discuss the relevance of these observations to the problem of infrared divergences that scattering amplitudes suffer from.
@@ IV Celestial Goldstone particles
$\begin{split}Q_{\lambda}&=\int_{\Sigma}d^{2}\vec{x}\,\lambda(\vec{x})\,\rho(\mathbf{x})\,,\\ Q_{T}&=\frac{2}{\kappa^{2}}\int_{\Sigma}d^{2}\vec{x}\,T(\vec{x})\,\mathcal{M}(\mathbf{x})\,,\end{split}$ (25)
where $\Sigma$ is a cut of $\mathscr{I}$, i.e., a surface of constant $u=u_{0}$. Here we pause to make contact with the literature. Recalling that $\mathscr{I}$ happens to be the future null conformal boundary of Minkowski spacetime, the current density $\rho(\mathbf{x})$ is identified with the pullback to $\mathscr{I}$ of the radial component of the electric field [12], while $\mathcal{M}(\mathbf{x})$ is identified with the BMS supermomentum as given in [24, 25] which can further be viewed as a component of the carrollian stress tensor [31, 39]. The charges (25) are the true canonical generators of asymptotic symmetries only when $\Sigma$ lies close to spatial infinity, i.e., when $u_{0}\to-\infty$. This is the region of spacetime where radiation is absent, and where the current densities are actually conserved. Away from spatial infinity, the current densities rather satisfy the classical evolution equations [15]
@@
On the right-hand sides of these equations $A_{i}(\mathbf{x})$ and $C_{ij}(\mathbf{x})$ (with $N_{ij}=\partial_{u}C_{ij}$) are the carrollian conformal fields carrying photons and gravitons, while $j_{u}$ and $T_{uu}$ are respectively the flux of electric charge and flux of energy carried away by matter fields. Equation (27) is called ‘Bondi mass loss formula’, as it expresses the fact that the total mass of the system contained within spacetime decreases as energy is carried away in the form of radiation. Going to the quantum theory, equations (26)-(27) should be understood as expectation value in some excited state $|\Psi\rangle$ made of ‘classical’ radiation (coherent superposition of massless particles),
$\langle\Psi|\partial_{u}\rho|\Psi\rangle=-\langle\Psi|\partial_{u}\partial^{i}A_{i}+e^{2}j_{u}|\Psi\rangle\,,$ (28)
and
$\begin{split}\langle\Psi|\partial_{u}\mathcal{M}|\Psi\rangle&=\frac{1}{4}\langle\Psi|\partial^{i}\partial^{j}N_{ij}+C^{ij}\partial_{u}N_{ij}|\Psi\rangle\\ &-\kappa^{2}\,\langle\Psi|T_{uu}|\Psi\rangle\,.\end{split}$ (29)
Note that the operators appearing on the right-hand side of these equations are composite operators made from the fundamental carrollian conformal fields (such as $C_{ij}(\mathbf{x})$). This contrasts with the interpretation given in [31] that they are background fields which induce non-conservation of the current densities $\rho(\mathbf{x})$ and $\mathcal{M}(\mathbf{x})$. Moreover, one can recognize equation (23) of [31] to be the Ward identity associated with unbroken asymptotic symmetries, while we proved here that they must be spontaneously broken. In vacuum $|\Psi\rangle=|0\rangle$, the right-hand sides of (28)-(29) vanish. For a quantum state $|\Psi\rangle$ made of localized excitations, away from the latter the currents are conserved, which we can equivalently write in the usual way as current conservation modulo contact terms (delta distributions of the form $\delta(\mathbf{x}-\mathbf{x}_{i})$), namely
$\begin{split}\langle\partial_{u}\rho(\mathbf{x})\,O_{1}(\mathbf{x}_{1})\,...\,O_{n}(\mathbf{x}_{n})\rangle&=contact\,,\\ \langle\partial_{u}\mathcal{M}(\mathbf{x})\,O_{1}(\mathbf{x}_{1})\,...\,O_{n}(\mathbf{x}_{n})\rangle&=contact\,.\end{split}$ (30)
The precise form of the contact terms will not be of importance in what follows.
@@
where we have assumed the vacuum to be time-translation invariant, i.e., $H|0\rangle=0$. Then carrollian current conservation implies, provided we analyse this equation away from the contact terms appearing in (30),
@@ V Discussion
where the remainder $R(\mathbf{x}_{0},\mathbf{x}_{1},\mathbf{x}_{2},...)$ accounts for the contributions of all other primaries and descendants. Provided we are able to determine the correlator $\langle\rho(\mathbf{x}_{0})G(\vec{x}_{2})...\,\rangle$, using (30) and (41) we can in principle infer the time-derivative $\partial_{u_{0}}R(\mathbf{x}_{0},\mathbf{x}_{1},\mathbf{x}_{2},...)$ of the remainder (away from coincident points). This potentially places interesting constraints on the carrollian operator product expansion, and further motivates the development of a carrollian conformal bootstrap as envisioned in [39].

### v3

@@ Abstract
The soft photon and soft graviton theorems of Weinberg are known to derive from conservation laws associated with asymptotic symmetries. Within the corresponding classical theories, one often speaks of spontaneous symmetry breaking and vacuum degeneracy, but a genuine quantum description of this phenomenon has largely been lacking. Here we establish spontaneous breaking of asymptotic symmetries and the existence of Goldstone ‘particles’ using exclusively the language of quantum field theory. This is made possible through the reformulation of massless scattering theory in terms of carrollian conformal field theory, and the observation that soft theorems correspond to Ward identities of broken symmetries. A suitable version of Goldstone theorem shows that there must exist zero-momentum particles described by conformal fields on the celestial sphere, in agreement with the common lore. More specifically, these belong to unitary representations in the discrete series of the Lorentz group, and are therefore naturally equipped with logarithmic two-point functions. We discuss the relevance of these observations to the problem of infrared divergences that scattering amplitudes suffer from.
@@ IV Celestial Goldstone particles
$\begin{split}Q_{\lambda}(u)&=\int_{\Sigma_{u}}d^{2}\vec{x}\,\lambda(\vec{x})\,\rho(\mathbf{x})\,,\\ Q_{T}(u)&=\frac{2}{\kappa^{2}}\int_{\Sigma_{u}}d^{2}\vec{x}\,T(\vec{x})\,\mathcal{M}(\mathbf{x})\,,\end{split}$ (25)
where $\Sigma_{u}$ is a cut of $\mathscr{I}$, i.e., a surface of fixed $u$. Here we pause to make contact with the literature. Recalling that $\mathscr{I}$ happens to be the future null conformal boundary of Minkowski spacetime, the current density $\rho(\mathbf{x})$ is identified with the pullback to $\mathscr{I}$ of the radial component of the electric field [12], while $\mathcal{M}(\mathbf{x})$ is identified with the BMS supermomentum as given in [24, 25] which can further be viewed as a component of the carrollian stress tensor [31, 39]. The charges (25) are the true canonical generators of asymptotic symmetries only when $\Sigma_{u}$ lies close to spatial infinity, i.e., when $u\to-\infty$. This is the region of spacetime where radiation is absent, and where the current densities are actually conserved. Away from spatial infinity, the current densities rather satisfy the classical evolution equations [15]
@@
On the right-hand sides of these equations $A_{i}(\mathbf{x})$ and $C_{ij}(\mathbf{x})$ (with $N_{ij}=\partial_{u}C_{ij}$) are the carrollian conformal fields carrying photons and gravitons, while $j_{u}$ and $T_{uu}$ are respectively the flux of electric charge and flux of energy carried away by matter fields. Equation (27) is called ‘Bondi mass loss formula’, as it expresses the fact that the total mass of the system contained within spacetime decreases as energy is carried away in the form of radiation. As emphasized previously, the charges $Q_{\lambda,T}(u)$ are canonical generators only in the limit $u\to-\infty$, precisely where the non-conservation terms on the right-hand side of (26)-(27) vanish. We will make use of this fact in what follows.
@@
where we have assumed the vacuum to be time-translation invariant, i.e., $H|0\rangle=0$. Then carrollian current conservation near $u\to-\infty$ implies
@@ V Discussion
where the remainder $R(\mathbf{x}_{0},\mathbf{x}_{1},\mathbf{x}_{2},...)$ accounts for the contributions of all other primaries and descendants. Provided we are able to determine the correlator $\langle\rho(\mathbf{x}_{0})G(\vec{x}_{2})...\,\rangle$, using (26) we can potentially infer the time-derivative $\partial_{u_{0}}R(\mathbf{x}_{0},\mathbf{x}_{1},\mathbf{x}_{2},...)$ of the remainder. This potentially places interesting constraints on the carrollian operator product expansion, and further motivates the development of a carrollian conformal bootstrap as envisioned in [39].